# Anomalous and nonlinear Hall effects in two-dimensional CrPS4

### v1

@@ Abstract
Few-layer CrPS4 is a two-dimensional (2D) magnetic material with excellent stability in ambient environment, which attracted significant interest in recent research. Here, via first-principles calculations, we show that 2D CrPS4 hosts a variety of anomalous Hall transport phenomena, owing to its layer-dependent magnetism and symmetry character. Monolayer CrPS4 can display a sizable linear anomalous Hall effect, while its nonlinear anomalous Hall response is forbidden. In contrast, bilayer CrPS4 can produce pronounced intrinsic nonlinear anomalous Hall response from Berry-connection polarizability, in the absence of linear anomalous Hall effect. We clarify that the large peaks in these responses originate from gapped Dirac points in the band structure. Furthermore, we show that linear anomalous Hall effect can be induced and controlled in bilayer CrPS4 by gate electric field or in-plane magnetic field, which break the spacetime inversion symmetry. Our findings unveil the interesting layer-dependent Hall transport physics in 2D CrPS4 magnets, suggesting its potential in electronic and spintronic device applications.
@@ I Introduction
The various Hall effects are of fundamental importance in condensed matter physics. A prominent example is the (linear) anomalous Hall effect (AHE), where a Hall voltage is induced by a longitudinal current flow in a magnetic material [1]. An important advance in the past two decades is the recognition that this effect contains an intrinsic contribution, i.e., an response determined solely by the material’s band structure, which manifests an intriguing band geometric property, the Berry curvature of electronic bands [2, 3]. AHE is a time-reversal-odd ($\mathcal{T}$-odd) effect, meaning that reversal of magnetic moments in the system (corresponding to a time-reversal operation) must flip the sign of the response coefficient. The $\mathcal{T}$-odd Hall effect was also discovered in nonlinear response of magnets, giving a Hall current quadratic in the driving $E$ field, i.e., $j_{\text{H}}\propto E^{2}$ [4, 5, 6]. Importantly, this nonlinear AHE also contains an intrinsic contribution, which is determined by another band geometric quantity, the Berry-connection polarizability (BCP) [6, 7], and can be further connected to the quantum metric of band structure [5, 8, 9]. Note that this effect is distinct from the $\mathcal{T}$-even nonlinear AHE that has been widely studied in nonmagnetic materials and does not have an intrinsic contribution [10, 11, 12, 13]. Recently, such intrinsic nonlinear AHE has been experimentally detected in a few magnets, such as few-layer MnBi2Te4 [14, 8] and Mn3Sn [15].
@@
Meanwhile, a focus of current materials research is to explore two-dimensional (2D) magnets [16, 17, 18]. Several material classes, like CrI3 [19], Fe3GeTe2 [20, 21], Cr2Ge2Te6 [22], and MnBi2Te4 [23, 24], have been realized. Due to the reduced dimensionality, a big advantage of 2D magnets is that their properties can be flexibly tuned, e.g., by gating, doping, strain, external fields, and hetero-structure engineering. Nevertheless, it was noted that most of 2D magnets studied so far are not stable in ambient environment, posing difficulties for applications. In this regard, the recently realized 2D magnet CrPS4 starts to attract increasing attention because of its excellent air stability [25]. It was shown that its magnetic ordering can persist down to monolayer limit [25]. The physical properties of CrPS4, such as spin dynamics, layer-dependent magnetism, optical response, electric and thermal transport properties, have been studied in several recent works [26, 27, 28, 29, 30, 25, 31, 32, 33, 34, 35, 36, 37]. As a semiconductor, the type and density of its charge carriers can be conveniently controlled by gating [35, 36]. Devices based on 2D CrPS4, such as field effect transistors and magnetic tunnel junctions, have been studied in experiment [36, 38, 37].
Here, we note that 2D CrPS4 actually offers a good platform to study a variety of Hall effects. First, 2D CrPS4 with odd number of layers is ferromagnetic (FM), which allows linear AHE. Second, even-layer CrPS4 has a compensated antiferromagnetic (AFM) ordering, which preserves spacetime inversion ($\mathcal{PT}$) symmetry. This forbids the linear AHE. However, the intrinsic nonlinear AHE is permitted. It should be noted that the symmetry condition of this nonlinear effect in 2D systems is a bit stringent: It requires the absence of any out-of-plane rotation axis [5, 6]. CrPS4 naturally satisfies this condition due to its monoclinic crystal structure. Third, although linear AHE is forbidden in the ground state of even-layer CrPS4, external fields may break the $\mathcal{PT}$ symmetry and generate a field-induced AHE.
@@
Motivated by these considerations, in this work, we perform a study of the various Hall responses in 2D CrPS4 via first-principles calculations. We focus on monolayer and bilayer CrPS4 as representatives. For monolayer CrPS4, we show that it has a sizable anomalous Hall conductivity in the case of hole doping. Its value can be strongly influenced by the magnetization direction. For bilayer CrPS4, we show that the intrinsic nonlinear AHE is determined by a single response coefficient, which can achieve a large value $\sim 1$ mA/V2 under hole doping. The origin of the large contribution in BCP is traced to a band anti-crossing region in valence bands. Furthermore, to induce linear AHE in bilayer CrPS4, we break the $\mathcal{PT}$ symmetry in two ways. The first is by a gate electric field (in the out-of-plane direction). The resulted net AHE response has a sign determined by the gate field direction, and the Hall current has a dominant distribution in one of the layers. The second way is to apply an in-plane magnetic field, which causes canting of the local moments and leads to an in-plane magnetization. We show that the resulting AHE can be quite large under moderate applied field strength. The two methods are also discussed from the perspective of layer Hall effect. Our results reveal interesting Hall transport physics in the 2D magnet CrPS4, which may find potential applications in novel electronic and spintronic devices.
@@ II Computation method
Figure 1: (a) Top view of a CrPS4 layer. The black rectangle marks the conventional cell. (b) Lattice structure of bulk CrPS4. It has a $A$-type AFM ground state. The local moments are in the out-of-plane direction, as indicated by the red arrows in the figure.
@@ III Structure and magnetic ordering
Bulk CrPS4 is a vdW layered material with monoclinic crystal structure belonging to the $C2/m$ space group (see Fig. 1). It was first synthesized and studied in the 1970s [48]. A recent experimental characterization reported that its lattice parameters (for the conventional cell) are $a=10.856\,\text{\AA}$, $b=7.247\,\text{\AA}$, and $c=6.135\,\text{\AA}$ [29]. As shown in Fig. 1(a), in a layer, each Cr cation is surrounded by six S anions, forming a CrS6 octahedron with slight Jahn-Teller distortion. These CrS6 octahedra form quasi-1D chains along $y$ direction in Fig. 1(a), which are further inter-connected by PS4 tetrahedra, resulting in a rectangle lattice in the 2D plane. Due to the relatively weak vdW interaction, a slight slip along a axis direction occurs between neighboring layers, causing angle $\beta=88.116^{\circ}$, which deviates slightly from $90^{\circ}$ (see Fig. 1(b)). The bulk structure of CrPS4 preserves the spatial inversion symmetry $\mathcal{P}$. However, due to the in-plane anisotropy as described above, there is no vertical rotation axis (only a twofold axis along $y$ exists).
The magnetic ground state of CrPS4 is well established in previous studies [25, 29]. The magnetism comes mainly from the Cr3+ ions with a magnetic moment about $3.0\mu_{B}$. The magnetic coupling within each layer is FM and the moments are in the out-of-plane direction; whereas the coupling between neighboring layers is AFM. This results in an overall $A$-type AFM ordering for bulk CrPS4, with Néel vector along $z$ axis, as shown in Fig. 1(b). The Néel temperature reported in experiment is about 35 to 38 K [28, 30, 29, 25, 38].
@@ IV Anomalous Hall response in monolayer
Due to $A$-type magnetic ordering, odd-layer CrPS4 has an uncompensated net magnetization. For monolayer CrPS4, this reduces to the FM ordering, with all magnetic moments in the out-of-plane direction ($z$ direction in our setup). The band structure for this ground state is plotted in Fig. 3(c) The result shows a semiconductor with a bandgap $\sim$1.20 eV, which agrees with previous studies [25].
@@
where $[d\bm{k}]$ is the short notation of $\sum_{n}d\bm{k}/(2\pi)^{2}$, $n$ is the band index, $f_{n\bm{k}}$ is the Fermi-Dirac distribution function, and
@@
In Fig. 4(a), we show the calculation result of $\sigma_{xy}$ as a function of the chemical potential $\mu$. One can observe several peaks in both hole-doped and electron-doped regions. The peak at $\mu\sim-0.32$ eV can reach up to $\sim-600$ S/cm, which is a quite large value. To understand the origin of this peak, we calculate the Berry curvature distribution in $k$ space summed over states below $-0.32$ eV. The result is shown in Fig. 4(b). One can see that large Berry curvature comes from two hot spots sitting on the $\Gamma$-$K$ path. These hot spots correspond to the small-gap regions indicated in Fig. 4(c). It should be noted that here, the two bands do not cross, instead, they form an anti-crossing with a small gap $\sim 0.014$ eV. Such anti-crossing may be considered as a 2D gapped Dirac point. From symmetry analysis, we construct the following $k\cdot p$ effective model for this gapped Dirac point:
@@
The magnetic anisotropy in 2D CrPS4 is not very large. It was shown in experiment that the magnetic moment direction can be rotated by relatively small applied magnetic field $\sim 90$ mT [25]. Hence, we also consider the AHE response in monolayer when the magnetization is rotated to an in-plane direction. In Fig. 4(d), we show the calculated $\sigma_{xy}$ for magnetization along the $x$ direction. Again, several peaks can be observed in the figure. The peak values are comparable to those in Fig. 4(a). The first peak in valence band has a value $\sim 370$ S/cm. Furthermore, for magnetization along $y$ direction, $\sigma_{xy}$ is identically zero. This is because in this configuration, the vertical mirror $\mathcal{M}_{y}$ symmetry is preserved, which forbids the AHE.
@@
Figure 5: (a) Lattice structure of bilayer CrPS4, which has $A$-type AFM ground state. (b) Calculated band structure for bilayer CrPS4 in AFM ground state.
@@
is the BCP tensor, and $\partial_{a}\equiv\partial_{k_{a}}$. Like Berry curvature, BCP is a gauge-invariant quantity. Physically, it characterizes the shift of a electron wavepacket enter by an applied electric field [4]. Near hot spots formed by a pair of bands, BCP can be further connected to the quantum metric of band structure [9].
@@
For an in-plane $E$ field in the direction $\bm{E}=E(\hat{x}\cos{\phi}+\hat{y}\sin{\phi})$, where $\phi$ is the polar angle measured from $x$ axis, the resulting Hall current $j_{\text{H}}$ in the direction normal to $E$ field can be expressed as
@@
Figure 6: (a) Calculated intrinsic nonlinear Hall conductivity $\chi_{yxx}$ as a function of chemical potential for bilayer AFM CrPS4. (b) $k$-space distribution of BCP dipole for states below $\mu=-0.3$ eV. The hot spots on $\Gamma$-$K$ path correspond to the two gapped Dirac points marked in the enlarged band structure in (c).
@@
where the wavevector $k$ is measured from the point, $\sigma$ and $\tau$ are two sets of Pauli matrices, $w$, $v$’s, $\alpha$, and $m$ are real model parameters. The last term in (10) opens a small gap $\sim 0.007$ eV between the two doubly degenerate bands. As shown in Ref. [6, 5], such gapped Dirac point tends to enhance BCP hence makes significant contribution to the nonlinear Hall response. In Fig. 6(c), there is another hot spot on the $k_{y}<0$ portion of $\Gamma$-$K$ path, with a comparatively weaker BCP dipole distribution. This corresponds to another gapped Dirac point at energy slightly below $-0.3$ eV (see Fig. 6(b)). The form of its effective model is the same as Eq. (10), since the local symmetry is the same.

### v2

# Tunable linear and nonlinear anomalous Hall transport in two-dimensional CrPS4

@@ Abstract
Few-layer CrPS4 is a two-dimensional (2D) magnetic material with excellent stability in ambient environment, which attracted significant interest in recent research. Here, via first-principles calculations, we show that 2D CrPS4 hosts a variety of intriguing anomalous Hall transport phenomena, owing to its layer-dependent magnetism and symmetry character. Monolayer CrPS4 can display a sizable linear anomalous Hall effect, while its nonlinear anomalous Hall response is forbidden. In contrast, bilayer CrPS4 can produce pronounced intrinsic nonlinear anomalous Hall response from Berry-connection polarizability, in the absence of linear anomalous Hall effect. We clarify that the large peaks in these responses originate from gapped Dirac points in the band structure. Furthermore, we show that bilayer CrPS4 also hosts a layer Hall effect, where the charge Hall current in the two layers are equal in magnitude but flow in opposite directions. This effect can be detected as an anomalous Hall signal induced by an applied gate field. Finally, even-layer CrPS4 also exhibits pronounced in-plane anomalous Hall effect, where the Hall signal is linearly proportional to the applied in-plane magnetic field. Our findings unveil the rich and interesting Hall transport phenomena in 2D CrPS4 magnets, suggesting its potential in electronic and spintronic device applications.
@@ I Introduction
The various Hall effects are of fundamental importance in materials science and device applications. A prominent example is the (linear) anomalous Hall effect (AHE), where a Hall voltage is induced by a longitudinal current flow in a magnetic material [1]. It offers an important characterization of magnetic materials and a convenient method for detecting magnetization dynamics. A key advance in the past two decades is the recognition that AHE contains an intrinsic contribution, i.e., an response determined solely by the material’s band structure, which manifests an intriguing band geometric property, namely, the Berry curvature of electronic bands [2, 3]. AHE is a time-reversal-odd ($\mathcal{T}$-odd) effect, meaning that reversal of magnetic moments in the system (corresponding to a time-reversal operation) must flip the sign of the response. The $\mathcal{T}$-odd Hall effect was also discovered in nonlinear response of magnets, giving a nonlinear Hall current quadratic in the driving $E$ field, i.e., $j_{\text{H}}\propto E^{2}$ [4, 5, 6]. Importantly, this nonlinear AHE also contains an intrinsic contribution, which is determined by another band geometric quantity, the Berry-connection polarizability (BCP) [6, 7], and it can be further connected to the quantum metric of band structure [5, 8, 9]. Note that because of its $\mathcal{T}$-odd nature, this intrinsic nonlinear AHE can occur only in magnetic materials; it is distinct from the $\mathcal{T}$-even nonlinear AHE that has been widely studied in nonmagnetic materials and does not have an intrinsic contribution [10, 11, 12, 13]. So far, the intrinsic nonlinear AHE has been experimentally detected in only a few materials, such as few-layer MnBi2Te4 [8, 14] and Mn3Sn [15]. However, these systems each has certain drawbacks that are not ideal for studying intrinsic nonlinear AHE. For example, MnBi2Te4 preserves a threefold rotational symmetry along $z$ direction, which forbids intrinsic nonlinear AHE in the $xy$-plane [5, 6], and some extrinsic mechanism, e.g., by using low-symmetry substrate or capping layers to break the $C_{3z}$ symmetry [8], is need for observing the intrinsic nonlinear AHE. In such a case, the detected response does not represent an intrinsic property of the magnet itself, but should be attributed to the interface or the whole combined heterostructure. For Mn3Sn, it also requires the strong interfacial effect at its interface with heavy-metal substrate [15]. In addition, Mn3Sn is not a van der Waals (vdW) layered material, making it more demanding in experimental fabrication of desired interfacial configurations. Therefore, at current stage, it is still an urgent task to find suitable magnetic materials that host intrinsic nonlinear AHE as its bulk property (i.e., without requiring interfacial effects). And it would be even better if such materials belong to the vdW 2D materials family.
@@
Indeed, 2D magnetic materials have been a focus of recent research [16, 17, 18]. Due to the reduced dimensionality, a big advantage of 2D magnets is that their properties can be flexibly tuned, e.g., by gating, doping, strain, external fields, and forming vdW heterostructures. Combined with magnetism, 2D magnetic materials are believed to hold tremendous potential for the next-generation information devices. Several material classes, like CrI3 [19], Fe3GeTe2 [20, 21], Cr2Ge2Te6 [22], and MnBi2Te4 [23, 24], have been realized. Nevertheless, it was noted that most of 2D magnets studied so far are not stable in ambient environment, posing difficulties for applications. In this regard, the recently realized 2D magnet CrPS4 starts to attract increasing attention because of its excellent air stability [25]. It was shown that its magnetic ordering can persist down to monolayer limit [25]. The fascinating properties of CrPS4, such as spin dynamics, layer-dependent magnetism, optical response, electric and thermal transport properties, have been extensively investigated in recent works [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. As a semiconductor, the type and density of its charge carriers can be conveniently controlled by gating [39, 40]. Devices based on 2D CrPS4, such as field effect transistors and magnetic tunnel junctions, have been studied in experiment [40, 41, 44]. However, the anomalous Hall transport properties, which is an important characteristic of magnetic materials, have not been studied in 2D CrPS4 yet.
@@
In this work, we reveal that 2D CrPS4 actually offers a good platform to study a variety of interesting Hall effects, especially the intrinsic nonlinear AHE. First, 2D CrPS4 with odd number of layers is ferromagnetic (FM), which allows linear AHE. Second, even-layer CrPS4 has a compensated antiferromagnetic (AFM) ordering, which has vanishing linear AHE but allows intrinsic nonlinear AHE. As mentioned above, the symmetry condition of intrinsic nonlinear AHE is stringent: It requires the absence of any out-of-plane rotation axis [5, 6]. CrPS4 stands out in this respect, as it naturally satisfies the condition due to its low-symmetry monoclinic crystal structure. Based on first-principles calculations and taking monolayer and bilayer CrPS4 as representatives, we quantitatively evaluate these intrinsic Hall response coefficients. For monolayer CrPS4, we show that it has a sizable anomalous Hall conductivity in the case of hole doping. Its value can be strongly tuned by the magnetization direction. For bilayer CrPS4, we show that the intrinsic nonlinear AHE is determined by a single response coefficient, which can achieve a large value $\sim 1$ mA/V2 under hole doping. The origin of the large contribution in BCP is traced to a band anti-crossing region in valence bands. Furthermore, we unveil that bilayer CrPS4 actually hosts a layer Hall effect, where the linear charge Hall current in the two layers are flowing in opposite directions. This effect can be probed as a net linear charge Hall signal induced by a gate electric field. The resulting current flow has a dominant distribution in one of the layers and the resulting Hall signal is odd in the applied gate field. Finally, we propose that a linear AHE signal can also be generated in even-layer CrPS4 by an in-plane magnetic field and the response is odd in the magnetic field, an effect known as the in-plane AHE. Based on reported experimental data, we show that the result can be quite large under moderate applied field strength. Our work reveals interesting Hall transport effects in the 2D air-stable magnet CrPS4, which form the basis for novel device applications in electronics and spintronics.
@@ II Computation method
Figure 1: (a) Top view of a CrPS4 layer. The black rectangle marks the conventional cell. (b) Lattice structure of bulk CrPS4. It has a AA-type AFM ground state. The local moments are in the out-of-plane direction, as indicated by the red arrows in the figure.
@@ III Structure and magnetic ordering
Bulk CrPS4 is a vdW layered material with monoclinic crystal structure (see Fig. 1). It was first synthesized and studied in the 1970s [59, 60, 61]. A recent experimental characterization reported that its lattice parameters (for the conventional cell) are $a=10.856\,\text{\AA }$, $b=7.247\,\text{\AA }$, and $c=6.135\,\text{\AA }$ [30]. As shown in Fig. 1(a), in a layer, each Cr cation is surrounded by six S anions, forming a CrS6 octahedron with slight Jahn-Teller distortion. These CrS6 octahedra form quasi-1D chains along $y$ direction in Fig. 1(a), which are further inter-connected by PS4 tetrahedra, resulting in a rectangle lattice in the 2D plane. Due to the relatively weak vdW interaction, a slight slip along a axis direction occurs between neighboring layers, causing angle $\beta=88.116^{\circ}$, which deviates slightly from $90^{\circ}$ (see Fig. 1(b))
It is noted that two initial experiments in 1977 reported slightly different crystal structures of bulk CrPS4  [59, 60]. Using X-ray diffraction method, Ref. [59] reported a crystal structure with $C2/m$ symmetry, i.e., with preserved inversion symmetry. On the other hand, Ref. [60] reported a very close structure but with weak inversion symmetry breaking, leading to a $C2$ symmetry. Comparing the two, the symmetry difference is associated with slight atomic position shift in each layer, mainly on the S site. Most subsequent research works on CrPS4 just quote the result, $C2/m$ [25, 34, 48, 62, 63, 64, 65, 66] or $C2$ [27, 29, 30, 31, 32, 37, 42, 43, 45, 49, 61, 67, 68, 69], from either one of Refs. [59, 60]. And in fact, because of the two structures are very close, they do not make any qualitative difference in explaining experimental results reported so far (within experimental accuracy). Our DFT calculations indicate that at least for 3D bulk CrPS4, monolayer CrPS4, and bilayer CrPS4, the $C2/m$ structure is energetically more stable than the $C2$ structure. Our test calculation also shows that the small structural difference makes little effect on the Hall transport properties, e.g., the linear AHE coefficient between the two are almost identical. Based on these considerations, we will take the $C2/m$ crystal structure for the investigation here. It follows that the bulk structure of CrPS4 preserves the spatial inversion symmetry $\mathcal{P}$ to a good extent. However, due to the in-plane anisotropy as described above, there is no vertical rotation axis (only a twofold axis along $y$ exists).
The magnetic ground state of CrPS4 has been well established in previous studies [25, 30]. The magnetism comes mainly from the Cr3+ ions with a magnetic moment about $3.0\mu_{B}$. The magnetic coupling within each layer is FM and the moments are in the out-of-plane direction; whereas the coupling between neighboring layers is AFM. This results in an overall $A$-type AFM ordering for bulk CrPS4, with Néel vector along $z$ axis, as shown in Fig. 1(b). The Néel temperature reported in experiment is about 35 to 38 K [25, 28, 30, 31, 41].
@@ IV Anomalous Hall response in monolayer
Due to the $A$-type magnetic ordering, odd-layer CrPS4 has an uncompensated net magnetization. For monolayer CrPS4, this reduces to the FM ordering, with all magnetic moments in the out-of-plane direction ($z$ direction in our setup). The band structure for this ground state is plotted in Fig. 3(c) The result shows a semiconductor with a bandgap $\sim$1.20 eV, which agrees with previous studies [25].
@@
where $d\bm{k}]$ is the short notation of $\sum_{n}d\bm{k}/(2\pi)^{2}$, $n$ is the band index, $f_{n\bm{k}}$ is the Fermi-Dirac distribution function, and
@@
In Fig. 4(a), we show the calculation result of $\sigma_{xy}$ as a function of the chemical potential $\mu$. One can observe several peaks in both hole-doped and electron-doped regions. The peak at $\mu\sim-0.32$ eV can reach up to $\sim-600$ S/cm, which is a quite large value. To understand the origin of this peak, we calculate the Berry curvature distribution in $k$ space summed over states below 0.32-0.32 eV. The result is shown in Fig. 4(b). One can see that large Berry curvature comes from two hot spots sitting on the $\Gamma$-$K$ path. These hot spots correspond to the small-gap regions indicated in Fig. 4(c). It should be noted that here, the two bands do not cross, instead, they form an anti-crossing with a small gap $\sim 0.014$ eV. Such anti-crossing may be considered as a 2D gapped Dirac point. From symmetry analysis, we construct the following $k\cdot p$ effective model for this gapped Dirac point:
@@
The magnetic anisotropy in 2D CrPS4 is not very large. From our calculations, the energy difference between out-of-plane and in-plane magnetic configurations is about 0.04 meV per Cr atom. It was shown in experiment that the magnetic moment direction can be rotated by relatively small applied magnetic field $\sim 90$ mT [25]. Hence, we also consider the AHE response in monolayer when the magnetization is rotated to an in-plane direction. In Fig. 4(d), we show the calculated $\sigma_{xy}$ for magnetization along the $x$ direction. Again, several peaks can be observed in the figure. The peak values are comparable to those in Fig. 4(a). The first peak in valence band has a value $\sim 370$ S/cm. Furthermore, for magnetization along $y$ direction, $\sigma_{xy}$ is identically zero. This is because in this configuration, the vertical mirror $\mathcal{M}_{y}$ symmetry is preserved, which forbids the linear AHE.
@@
Figure 5: (a) Lattice structure of bilayer CrPS4, which has AA-type AFM ground state. (b) Calculated band structure for bilayer CrPS4 in AFM ground state.
@@
is the BCP tensor, and $\partial_{a}\equiv\partial_{k_{a}}$. Like Berry curvature, BCP is a gauge-invariant quantity. Physically, it characterizes the shift of a electron wavepacket enter by an applied electric field [4]. Near hot spots formed by a pair of bands, BCP can be further connected to the quantum metric of band structure [9]. From Eqs. (5, 6, 7), one can clearly see that the intrinsic response is purely Hall and dissipationless, i.e., $\chi_{abc}$ is antisymmetric in its first two indices [71]. As a consequence, the intrinsic conductivity $\chi_{abc}$ is also forbidden by any rotational symmetry normal to the transport plane. These key features and the theory have been successfully verified in experiments on MnBi2Te4 [8] and Mn3Sn [15]. Meanwhile, responses from extrinsic mechanisms do not share such characters [72, 73, 14]. This theory has found successful applications in explaining key features in experiments on MnBi2Te4 [8] and Mn3Sn [15].
@@
For an in-plane $E$ field in the direction $\bm{E}=E(\hat{x}\cos{\phi}+\hat{y}\sin{\phi})$, where $\phi$ is the polar angle measured from $x$ axis(as shown in Fig. 6(d)), the resulting Hall current $j_{\text{H}}$ in the direction normal to $E$ field can be expressed as
@@
Figure 6: (a) Calculated intrinsic nonlinear Hall conductivity $\chi_{yxx}$ as a function of chemical potential for bilayer AFM CrPS4. (b) $k$-space distribution of BCP dipole for states below $\mu=-0.3$ eV. The hot spots on $\Gamma$-$K$ path correspond to the two gapped Dirac points marked in the enlarged band structure in (c).(d) Sketch of nonlinear Hall current induced by $E$ field.
@@
where the wavevector $k$ is measured from the point, $\sigma$ and $\tau$ are two sets of Pauli matrices, $\tau^{\prime}s$ denote the two doubly degenerate anti-crossing bands, and $\sigma^{\prime}s$ denote the spin degree of freedom within each doubly degenerate band, $w$, $v$’s, $\alpha$, and $m$ are real model parameters. The last term in (10) opens a small gap $\sim 0.007$ eV between the two doubly degenerate bands. As shown in Ref. [5, 6], such gapped Dirac point tends to enhance BCP hence makes significant contribution to the nonlinear Hall response. In Fig. 6(c), there is another hot spot on the $k_{y}<0$ portion of $\Gamma$-$K$ path, with a comparatively weaker BCP dipole distribution. This corresponds to another gapped Dirac point at energy slightly below 0.3-0.3 eV (see Fig. 6(b)). The form of its effective model is the same as Eq. (10), since the local symmetry is the same.
The above discussion is for the AFM bilayer with top layer spin-down and bottom layer spin-up. As for the other energy-degenerate configuration with top layer spin-up and bottom layer spin-down, the nonlinear Hall response will flip sign. This is because the two magnetic configurations are connected by time reversal operation, and the BCP dipole induced nonlinear Hall response is a $\mathcal{T}$-odd effect.